OpenStudy (bigbosssaint21):

Identify the factors of x^2 + 49y^2. (x + 7y)(x + 7y) Prime (x - 7y)(x - 7y) (x + 7y)(x - 7y)

OpenStudy (neer2890):

what's 2nd option?

OpenStudy (bigbosssaint21):

that its prime

OpenStudy (neer2890):

yup, it's prime

OpenStudy (neer2890):

because (x+7y)(x+7y)=(x+7y)^2=x^2+49y^2+14xy which is not equal to x^2+49y^2 (x-7y)(x-7y)=(x-7y)^2=x^2+49y^2-14xy (x-7y)(x+7y)=x^2-(7y)^2=x^2-49y^2 because \[a ^{2}-b ^{2}=(a+b)(a-b)\]